A man has 11 pairs of socks in his sock drawer: 3 brown, 6 black, and 2 blue.
a): if the man reaches into his drawer with his eyes closed what is the fractional probability that the pair of socks he chooses is black?

Answer by jim_thompson5910(35256)

a)
P(2 black socks) = P(Black Sock AND Black Sock)

P(2 black socks) = P(Black Sock)*P(Black Sock)

P(2 black socks) = (6/11)*(5/10)

P(2 black socks) = 30/110

P(2 black socks) = 3/11


So the answer is 3/11
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------